OpenStudy (anonymous):

SIMPLE MATRIX QUESTION PLEASE HELP!! are the given vectors normal? a = (5,-2) and b = (6,15)

OpenStudy (anonymous):

@GEERKY42

OpenStudy (anonymous):

@sithsandgiggles

OpenStudy (anonymous):

@jagr2713

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@greg_d

OpenStudy (anonymous):

is the question asking if they are perpendicular to each other?

OpenStudy (anonymous):

@greg_d all it is asking is if it is "normal"

OpenStudy (anonymous):

I can't determine what or why or how it would be considered normal.. the rest of the homework was working on matrices

OpenStudy (anonymous):

if the question is, are they "normal" (as perpendicular) to each other, you can use the dot product...

OpenStudy (anonymous):

This is honestly the last problem and i don't know how i could do that..

OpenStudy (anonymous):

@greg_d

OpenStudy (anonymous):

ok, IF that is the question, we need to multiply them using the dot product. \[a.b=a_xb_x+a_yb_y\] since\[a.b=|a||b|cos(\theta) \] with theta beign the angle between the vectors. if that product is 0, said angle is 90º and they are indeed "normal" to each other

OpenStudy (anonymous):

could you possibly solve that for me so i can use it next time i see something like this?

OpenStudy (anonymous):

@greg_d

jimthompson5910 (jim_thompson5910):

Similar Example: Let's say you had the two vectors u = <1,2> v = <5,7> The dot product of u and v is u dot v = 1*5 + 2*7 = 5 + 14 = 19 Since the dot product is not 0, this means that u and v are not perpendicular

jimthompson5910 (jim_thompson5910):

In general u = <a,b> v = <c,d> u dot v = a*c + b*d u,v are vectors a,b,c,d are scalars

OpenStudy (anonymous):

@jim_thompson5910 @greg_d Thank you!!

OpenStudy (anonymous):

:D that examples will lead you to the answer !